OpenStudy (anonymous):

Factor: 3x squared minus 21x minus 54

OpenStudy (anonymous):

So the question looks like \[3x ^{2}-21x-54\]? Correct?

OpenStudy (anonymous):

Yep

OpenStudy (triciaal):

welcome

OpenStudy (anonymous):

Okay. What is one thing that you can factor out of all 3 parts of the equation?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Yep. So what are you left with once you factor that out?

OpenStudy (anonymous):

X-7x-18

OpenStudy (anonymous):

Yep. Does that factor?

OpenStudy (anonymous):

Like my teacher wants me to factor it like. for example x squared minus 7x minus 18. like the factored would be (×-9)(×+2)

OpenStudy (anonymous):

\[\frac{1}{3} \left(3 x^2-21 x-54\right)=x^2-7 x-18 \]

OpenStudy (triciaal):

**x^2 -7x-18 the first term is x^2

OpenStudy (anonymous):

Yep!! So you're answers would be?

OpenStudy (anonymous):

Thats what im asking it has a coefficiant and changes it completely how to do it

OpenStudy (anonymous):

like instead of the first term having a coefficient of one it has one of 3 and changes how to do the whole distributive property on it

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @bryce0706 X-7x-18 \(\color{blue}{\text{End of Quote}}\) 3 is common factor so u r left with (x^2 -7x -18) find two number if you multiply them you should product of AC and if you add or subtract them you should get middle term

OpenStudy (anonymous):

hi

OpenStudy (triciaal):

@bryce0706 do you need more help?